MODS1 Commissioning Run 5: UTC 2011 February 16

Observer(s): R. Pogge & J. Antognini (OSU), M. Pedani (LBTO)
Telescope Operator: D. Gonzalez-Huerta (LBTO)
AO Support: A. Rakich, J. Hill & D. Miller remote from Tucson
Telescope Support: J. Little (LBTO)
Instrument Support: J. Morris (LBTO)

Summary

Clear most of the night but windy. We lost the last 4 hours of the night to high winds. Got some work done in between, mostly hampered by poor seeing.

Active optics tests undertaken from Tucson gathered data that should help tune up the active optics corrections, especially for off-axis stars which will be the rule for multi-object spectroscopy with this instrument.

Got our first taste of laser-cut multi-object mask designed using the prototype mms software. Also got a spectrophotometric wide long-slit mask and took a few spectral sequences before the weather went south.

Preparations

Caught up on comparison lamp spectra for all modes, as the telescope was sent to horizon at dawn for work that began immediately after we went to bed. With it locked at zenith this afternoon we shot prism and grating mode comparison lamp in dual, red-, and blue-only modes.

New slit masks arrived from Tucson, cut with the laser machine from masks designed with the prototype "mms" MODS Mask Software package (basically, lms for MODS). 4 masks installed:
  • LS60x5 - 5-arcsec wide spectrophotometric calibration short slit
  • StoneI_MOS - ID525427 MOS Test Mask
  • StoneF_MOS - ID532606 MOS Test Mask
  • Abell1689 - ID594500 MOS Field Mask

LS60x5 replaced the pinhole slit mask, Stone I, Stone F, and Abell1689 went into 18, 20, and 22, and the BruceMask in position 23 was temporarily removed to make room for the test masks for this run and likely the next. We are still using every-other packing because of the misaligned grabber.

Bug fixes in the lbttcs agent:
  • if AIRMASS>99.99, sets it to 99.99. This is the placeholder max value for when the telescope is near the horizon (e.g., requesting bias frames when horizon-pointing would result in huge values of AIRMASS in the headers).
  • If Dec was given without a sign (see yesterday's report), a weird value was sent to IIF. This was traced to the lbttcs agent which was incorrectly dealing with the sign-less (implicit +) case for dec entry for target and guide object coordinates. A better algorithm was implemented.

Details

All times are UTC unless otherwise noted.

0100 - A few isolated clouds at sunset - probably will not be photometric. These are likely the first outriders of the big low brewing off the Pacific.

0130 - While waiting for the sky to get dark, taking some MOS test mask calibrations. These will need some work... Cloud bits blowing past fast, including quite a few contrails.

0150 - Getting ready to open - First pointing preset... ACT0193
   IE = -67
   CA = +89

0154 - On-axis active preset to BS9108. Starting way out of focus, Andrew dialed in some initial correction and then we let the active optics take over. Started converging relatively quickly. Part way Andrew rolled out the global offset and let actitve optics continue.

Quite a bit of dome seeing in the WF images, we'll do some other startup and initial imaging while things settle down. They'll proabably revisit the testing around 0300 UTC. Seeing is around 1-arcsec.

0215 - Acquiring G191B2B, testing the new 60x5-arcsec spectrophotometric slit (LS60x5). We'll acquire 1Kx1K, then put in dual grating mode and snap a spectrum and see how it looks. Conditions are not perfectly photometric, but not bad either (clouds are discrete, and we have bright enough moon to see them on the all-sky camera).

First, however, the preset has to succeed... which it did. Seeing is ~1arcsec or less, collimation starting to converge

This slit is large enough, we can probably acquire directly through the slit in most cases w/o a field image.

0225 - First snap thru slit.

0235 - Messing around, writing a script to automate the offset calc better (crude prototype).

Dual Grating Mode, 180s per side. Seeing ~0.8 arcsec FWHM, RMS WF error ~180nm. Still dome seeing

Dual Prism followed - this is a nice absolute calibration slit.

Active Optics Tuneup

0300 -

Back to Observing

0435 - Checking pointing again after all the experiments of the past 1.5 hours.
   IE = -77
   CA = 95

0440 - Collimation Check - on-axis preset to BS9126

Preset failed with a GCS timeout waiting for probe motion. I see the request and response in the MODS AGw server log - network glitch of some kind not getting the message to GCS?

Resent the preset and it went right to the source.

First iteration on star, WF err is 1300nm, starting to converge subsequently. Seeing ~0.8-arcsec FWHM and about 130nm RMS WF error

Hiltner 600 - big slit observations

0448 - Active preset with a near-by star. Using LS60x5 specphot slit.

0452 - Slit acquisition image - should be already in the big slit... wow - right smack in the slit, no offset required. Our first

0455 - Dual Grating Mode -

PG0823+546 - SpecPhot star

0532 - On-axis preset to update collimation

0540 - Off-axis preset w/guide star - seeing 0.9 arcsec, WF error 100nm rms

0543 - Slit acquisition - using the XVista acqslit script to semi-automate just for this run

0548 - Running calibrations using a script for all 6 modes

Red IMCS took >60s to converge. Increasing to 120s inside the script and re-executing. We need to revisit the internal IMCS look-up tables for the mode presets, we seem to be off.

Ran the script, mostly done, while demoing something accidentally aborted the last exposure. My bad, but no big loss, this was mostly an "observing process" task, and so oops is part of the testing.

Stone F Field

0633 - Moving to the Stone F astrometric reference field. Coordinates uploaded using a skeleton .acq script, but without any acquisition command blocks (preset launched by hand from the MODS control panel GUI). Active preset is to an off-axis star.

Staged the StoneF_MOS mask as part of the process.

Seeing is 1.1-1.2 arcsec and a little swimmy, we're facing into a wind of up to 16 m/s. We'll do what we can.

0643 - First acquisition shot, no mask.

A bit of a scare, on extracting the mask, it showed MASKPOS=Empty on return. This was distressingly like what it reported back last September when we had a mask fall out during the inaugural run. We moved to zenith, locked the drives and rotator, and went up to inspect the mask system.

The mask was in the right place and not out of position. However the "flapper" with the mask in-stow-position sensor was a little ajar and loose - it is usually held in position by a combination of a snap hinge at the base of the flapper and a magnetic tab at the top. I reseated it and had David read back the "slitmask" status. This time it showed MASKPOS=STOW as it should. I wiggled it and worked it some, every now and then chekcing status (including querying status with the flapper in the mask load/remove position - it read "empty" as it should in that state).

We then did a couple of insert/retract cycles with telescope at Zenith, no further repeat of the sensor fault.

When I installed the new masks earlier today, the scissors lift was further from the instrument, and I had to reach in more to put the flapper back in place. I wonder if it was not slightly ajar and the work tonight was such that the magnetic tab disengaged and it sprang back a little. We will review this problem back in Columbus with the MODS2 mask system in the lab. There is no danger (except to possibly give the instrument PI a heart attack), but we'll keep an eye on it.

0730 - resuming work, back to the field. Wind is up, so David is rotating azimuth around to a position where he can check the wind speed w/o the building shadow. At least 1 gust up to 19m/s. Temp is 3°C, humidity 22%, wind out of the SW. After a while it appears that was just a rogue gust.

0741 - back to the target. We are at elevation 45° and now the sensor is reading EMPTY. Hmm. I translated the mask cassette to occupied positions 16 and 14 (next up), and they show STOW (i.e., occupied). Back to 18 shows empty. I wonder what is up with that? PARANGLE=40°

Going back to zenith... watching the sensor as we go. It asserted STOW at around El=80°. Step down in elevation by 5°
    90 - STOW
    85 - STOW
    80 - STOW
    75 - STOW
    70 - STOW
    65 - EMPTY at around 68
    60 - EMPTY
Now at El=60°, slewing rotator from 248° to 68° Transitioned to STOW at about 130°

So, perhaps what we are seeing is a slight flexure of the mask relative to the flapper sensor moving the outside surface of the mask frame (the "load side") in and out of sensor's trigger range. The trigger error in these proximity sensors is about +/-10-microns according to the manufacturer.

Stone I Field

0800 - Since the Stone I mask does not have this oddball problem, we'll go that target next. During the preset I watched the mask in-stow-position sensor and it showed no change.

However, once on target, the seeing is bad 1.5 to 2.8-arcsec in puffs. Wind is up, you can feel it through the floor, and we're facing about 90° to the wind. Gusts up to 16-19m/s

Am going to a nearby star as an on-axis active preset to check collimation, we started way out and the star for this field was off-axis enough to have a partially-vignetted WFS image.

Converged at only 300-400nm RMS. Seeing is puffing up now and then, gusts up to 17m/s during the period.

0820 - back to the target. Converged to 150nm RMS, but seeing is 1.5 to worse arcsec. Wind variable, 3 to 17m/s in the past 10 minutes.

0830 - watching the guide star bounce around and puff up to >2-arcsec.

0840 - Wind exceeding limits, closing up.

1200 - Winds still to strong to open, so we call it a night.

-- RichardPogge - 15 Feb 2011
Topic revision: r8 - 16 Aug 2011, RichardPogge
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback